OpenStudy (anonymous):

Eliminate the parameter and write a rectangular equation for x=t^2+3 [right below first equation] y=2t ??

OpenStudy (anonymous):

solve for t. so you have y=2t, divide both sides by 2 and you get t=y/2. Then plug this back into the first equation. so we will have x=(y/2)^2+3 which is x=y^2/4+3! :)

OpenStudy (anonymous):

Thank you, again! :)

OpenStudy (anonymous):

no problem :)